# Object Re-identification via Spatial-temporal Fusion Networks and Causal Identity Matching

### v1

@@ III-B Fusion Network
The baseline for the proposed ReID framework can be any appearance-based ReID method, which estimates appearance similarities between images. Object images from each camera pair $(c_{i},c_{j})$ are represented as $\mathbf{I}^{m}_{i},\mathbf{I}^{n}_{j}$, where $m$ and $n$ are the image indexes. Then, the appearance-based ReID methods estimate the visual similarity between two images as $S_{A}(\mathbf{I}^{m}_{i},\mathbf{I}^{n}_{j})$ that lies on $\left[0,1\right]$. The proposed framework is not reliant on typical appearance-based models.
To perform spatial-temporal ReID, Cho et al. [7] used only camera network topology to limit the gallery search range, effectively reducing the complexity of ReID. However, the spatial-temporal probability does not affect the final similarity. In contrast, previous studies [10, 33, 9] have merged both probabilities (i.e., appearance and spatial-temporal) with the same importance to obtain the joint probability. However, they neglected two key points. First, the domains of each probability are not the same. Second, both appearance and spatial-temporal probabilities can be imperfect. Hence, simply merging these probabilities is unreasonable.
In this study, we optimally combined visual similarities $S_{A}(\mathbf{I}^{m}_{i},\mathbf{I}^{n}_{j})$ and estimated spatial-temporal probability distributions $p_{ij}\left(\tau\right)$ through the fusion network named FusionNet. An input vector of the network for two images $(\mathbf{I}^{m}_{i},\mathbf{I}^{n}_{j})$ in the camera pair $(i,j)$ can be represented by
@@
where $S_{A}$ is an appearance similarity, and $S_{T}$ is a spatial-temporal vector. The $S_{T}$ vector between the images is defined by
@@
the notation $\tau^{mn}_{ij}$ represents the time difference between two images $\mathbf{I}^{m}_{i}$ and $\mathbf{I}^{n}_{j}$. The parameter $W$ denotes the size of a time window. The range of the $S_{T}$ vector depends on the value of $W$ and is determined around distributions of $p_{ij}(\tau^{mn}_{ij})$. For example, when $W=0$, $S_{T}$ becomes a scalar value given by $S_{T}(\mathbf{I}^{m}_{i},\mathbf{I}^{n}_{j})=p_{ij}(\tau^{mn}_{ij})$. If $W>0$, the $S_{T}$ vector has a dimension of $12W+$ 1. The value of $W$ can be adjusted to determine the amount of spatial-temporal information to be provided to the FusionNet. Then, the dimension of the input vector $\mathbf{x}^{mn}_{ij}$ for the FusionNet is $22W+$ 2.
(a) One-to-all re-identification without causality
(b) ID-to-ID re-identification considering causality
Figure 4: Comparisons of re-identification methodologies. Circles and boxes denote cameras and appearances, respectively. The number in the box denotes the ID of the appearance. Boxes filled with olive color are true positives with the query. Boxes with a dotted line indicate excluded appearances from the gallery. (a) One-to-all ReID using only a single appearance for the query performs many redundant and duplicated comparisons. It even compares different objects within the same camera.
@@
The FusionNet was designed based on a simple multi-layer perception. We found empirically that the FusionNet does not require a sophisticated deep neural network structure to estimate the final similarity. The network has one hidden layer with several nodes and a one-dimensional output layer, as illustrated in Fig. 2(a). For the activation function, we used the Rectified Linear Unit (ReLU) for nodes in the hidden layer and the sigmoid function for the output node. Then, the final output of the FusionNet, $S_{F}(\mathbf{I}^{m}_{i},\mathbf{I}^{n}_{j})$, is within the range $\left[0,1\right]$. To train the network, we optimized the binary cross-entropy loss as defined by
@@ IV Re-identification in
Previous studies have evaluated their methods using a fixed and refined gallery set. They selected a query image from the gallery and compared it with all other images in the gallery, which is a one-to-all single-shot comparison. This evaluation process is repeated until all queries are compared, and then all the results are averaged to compute the final ReID performance. Note that the gallery set contains multiple appearances for a single identity, leading to many duplicated comparisons as shown in Fig. 4a. These laboratory evaluation settings are not suitable for real-world scenarios at all, since multiple appearances of the object do not need to be evaluated individually.
In real-world surveillance scenarios, we have to consider additional issues beyond the laboratory settings. First, each identity can have multiple appearances in the video, and their appearances should be evaluated in ID-to-ID manner as explained in Fig. 4b. Second, the query and gallery sets are not fixed throughout all the test processes; instead, they can change according to the target objects and matching causality. Considering the issues in real-world scenarios, we additionally propose causal identity matching (CIM) from our previous work [12]. CIM includes three different methods (Section. IV-A– IV-C) for effective ReID in real-world scenarios. The overall framework of CIM is illustrated in Fig. 2(b).
@@ IV-A Appearance management for each object
As mentioned above, each object has multiple appearances along with its trajectory in the sequence. Assuming that a camera captures multiple frames per second, we can collect various images, which is advantageous for ReID. However, this also requires a significant amount of computation. We observed that many of the appearances are unreliable and redundant in many cases. For example, when another object occludes an object, the appearances of different objects can become very similar, resulting in unreliable appearances. Additionally, when an object remains in the same location and pose for a long period, multiple appearances of the object are identical, leading to redundant appearances. We here define the bounding box of the $m$-th object image $\mathbf{I}^{m}$ as $\text{bbox}(\mathbf{I}^{m})=\left[x^{m},y^{m},w^{m},h^{m}\right]$. We omit camera indexes $i,j$ for convenience at this step. In the video, the $m$-th object contains a set of appearances as
@@
where $N_{m}$ is the total number of the appearances in $m$-th object.
@@
Figure 5: Features distributions of objects. Example image patches reflect actual sizes of them.
Inspired by [34], we perform appearance management for each object. To detect occluded unreliable objects, the occlusion ratio between two different objects is calculated based on Intersection over Union (IoU) by $\text{IoU}\left(\mathbf{I}^{m},\mathbf{I}^{n}\right)=\frac{\text{bbox}(\mathbf{I}^{m})\cap\text{bbox}(\mathbf{I}^{n})}{\text{bbox}(\mathbf{I}^{m})\cup\text{bbox}(\mathbf{I}^{n})}$. In general, surveillance cameras are installed at a high position from the ground, looking down at an angle. This means that an overlapping object with a larger $y$-value is occluding another object. Therefore, we designed the occlusion ratio of $m$-th object by
@@
where $m\neq n$. When there are multiple bounding boxes occluding $\mathbf{I}^{m}$, the largest value among them will be used as the final occlusion ratio.
Next, we measure a self-overlapping ratio to prevent duplicate appearances from being included in the set $\mathcal{A}^{m}$. For simplicity, let $\mathbf{\hat{I}}^{m}\in\mathcal{A}^{m}$ is the last appearance in the set $\mathcal{A}^{m}$. Then, the self-overlapping ratio calculated by $\text{IoU}\left(\mathbf{I}^{m},\mathbf{\hat{I}}^{m}\right)$. Finally, we merge two ratios to define the appearance unreliability score of $\mathbf{I}^{m}$ by
@@
Figure 5 illustrates the feature distributions of four different objects. Without the proposed appearance management, they are not clearly separated in the feature space due to occlusions. Furthermore, each object has similar and redundant appearances when standing at the traffic signal. After performing appearance management, the unreliable appearances were appropriately removed, preserving the original distributions with a much smaller number of appearances.
@@ IV-B Top-$k$ Multi-shot Matching
After appearance management, objects contain more reliable appearances in their appearance sets. In order to compare two different appearance sets, many traditional multi-shot matching ReID methods have tried minimum (max) and average (avg.) aggregation approaches in general. However, the ‘max’ approach, which selects only the maximum similarity value among multiple appearance comparison results, can result in highly biased final outcomes. On the other hand, the ‘avg.’ approach, which averages all the comparison results, can result in overly smoothed final outcomes.
To mitigate these risks, we perform Top-$k$ multi-shot matching to compare two different appearance sets. Given two different appearance sets $\mathcal{A}^{m},\mathcal{A}^{n}$ with their cardinalities $|\mathcal{A}^{m}|=N_{m}$, $|\mathcal{A}^{n}|=N_{n}$, there are $N_{m}\times N_{n}$ appearance comparisons. Based on the FusionNet in Section III-B, the appearance similarities between the two sets are represented by
@@
Sort $S_{F}\left(\mathcal{A}^{m},\mathcal{A}^{n}\right)$ in descending order and then we get
@@
Finally, top-$k$ multi-shot matching is calculated by
@@
By adjusting $k$, we can determine how many results to utilize in the multi-shot appearance comparison. We empirically set the value of $k$ around $5$, which improved ReID performances as in Table. VII. The proposed top-$k$ approach is quite simple but effective. It does not require complex structures for handling multi-shot comparisons like video-based ReID methods [35, 36] but can utilize single-shot baselines. In addition, it implements other approaches by setting $k=1$ for the ‘max’ approach and $k=N_{m}\times N_{n}$ for the ‘avg.’ approach.
@@ IV-C Causal Identity Matching (CIM)
In Section III-A, camera network topology was built as $\mathbf{G}=(\mathbf{V},\mathbf{E})$, and FusionNet was trained based on this topology. The topology includes transition time distributions $p_{ij}$ between all possible camera pairs, regardless of their connection strengths, enabling FusionNet to learn these strengths for determining the similarity $S_{F}$. In real-world scenarios, we focus more on the target object (i.e., query) and its causality in the camera networks $\mathbf{G}$. The causality indicates where the target object is currently and where it is likely to move next based on the topology. In this setting, the query and gallery sets should not be fixed or pre-defined but should dynamically change according to the ReID results at each step.
To consider aforementioned conditions and perform the causal identity matching (CIM), we first define camera adjacent matrix by $\mathbf{C}\in\{c_{ij}|1\leq i\leq N_{cam},1\leq j\leq N_{cam}\}$, where $i,j$ are camera indexes, and $c_{ij}$ denotes a binary adjacent flag between cameras. When cameras $c_{i}$ and $c_{j}$ are adjacent, the flag is set to $c_{ij}=1$; otherwise, it is set to $c_{ij}=0$. We set all the diagonal terms in $\mathbf{C}$ to $c_{ii}=0$. Given multiple object paths, we can simply find adjacent cameras and overall connections between them. For example, assuming that there are multiple object paths such as $\left(c_{1}\rightarrow c_{2}\rightarrow c_{3}\right)$, $\left(c_{2}\rightarrow c_{3}\right)$, $\left(c_{2}\rightarrow c_{4}\right)$, and $\left(c_{4}\rightarrow c_{2}\right)$; then the camera adjacent matrix is defined by
@@
Note that the matrix can be asymmetric when considering a bidirectional camera network topology. More reliable adjacent connections can be estimated by voting on sufficient numbers of paths. In addition, unnecessary paths can be further removed through post-processing algorithms like transitive reduction [37] [^1] When there is a new path $\left(c_{1}\rightarrow c_{3}\right)$, it does not assign $c_{13}=1$, immediately. Instead, the algorithm tries to infer if there is an intermediate vertex $c_{2}$ between $c_{1}$ and $c_{3}$ from numerous training paths.. Excluding redundant paths helps to improve the overall efficiency of ReID.
Next, we assign dynamic gallery sets for the target object. Imagine that there is a target object (query) in the camera $c_{1}$. Then, the gallery set of the query comes from the cameras adjacent to $c_{1}$ by finding $(c_{1j}=1)$ in $\mathbf{C}$. Furthermore, the objects in the gallery are confirmed by referring to the corresponding transition time distributions $p_{1j}\left(t\right)$. Only the objects in the camera $j$ within the appearing time $t$ that satisfy $p_{1j}\left(t\right)>0$ are included in the gallery. Through this process, we can causally assign the gallery set using $\mathbf{C}$ and reduce many unnecessary matching candidates based on $p_{ij}\left(t\right)$.
An object in the gallery with the highest similarity to the query is considered a successful match, and the query is merged with the re-identified object. This increases the diversity of the query’s appearance, making the next step of ReID much more advantageous. In other words, although a single camera can only capture fixed poses of the object, progressive query merging easily overcomes this limitation. In this way, the updated query is used to progressively perform ReID on the next adjacent cameras, and the gallery is adaptively determined by $\mathbf{C}$ and $p_{ij}$ as well.
V Datasets
Index Cam 1 Cam 2 Cam 3 Cam 4 Cam 5 Cam 6 Cam 7 Cam 8 Cam 9 Cam 10 Cam 11 Total
# ID 456 240 494 570 696 430 667 788 159 843 408 2,038
# Annotated box 4,905 2,667 5,200 3,524 3,603 2,466 3,510 3,975 1,077 6,509 2,905 40,341
# Frames 17,979 17,973 17,986 14,300 17,997 17,958 17,991 17,961 17,987 17,963 17,883 -
Duration 19m 58s 19m 58s 19m 59s 15m 53s 19m 59s 19m 57s 19m 59s 19m 57s 19m 59s 19m 57s 19m 52s -
Color (%) [Uncaptioned image] 40.13 [Uncaptioned image] 38.75 [Uncaptioned image] 39.27 [Uncaptioned image] 39.47 [Uncaptioned image] 38.22 [Uncaptioned image] 41.16 [Uncaptioned image] 41.53 [Uncaptioned image] 37.69 [Uncaptioned image] 39.62 [Uncaptioned image] 37.01 [Uncaptioned image] 41.67 [Uncaptioned image] 38.96
[Uncaptioned image] 30.92 [Uncaptioned image] 30.42 [Uncaptioned image] 30.77 [Uncaptioned image] 35.79 [Uncaptioned image] 34.63 [Uncaptioned image] 30.23 [Uncaptioned image] 31.18 [Uncaptioned image] 36.42 [Uncaptioned image] 39.62 [Uncaptioned image] 36.18 [Uncaptioned image] 38.73 [Uncaptioned image] 34.40
[Uncaptioned image] 22.81 [Uncaptioned image] 24.58 [Uncaptioned image] 24.09 [Uncaptioned image] 20.70 [Uncaptioned image] 21.70 [Uncaptioned image] 21.86 [Uncaptioned image] 21.14 [Uncaptioned image] 20.30 [Uncaptioned image] 15.09 [Uncaptioned image] 21.59 [Uncaptioned image] 15.93 [Uncaptioned image] 20.85
[Uncaptioned image] 6.14 [Uncaptioned image] 6.25 [Uncaptioned image] 5.87 [Uncaptioned image] 4.04 [Uncaptioned image] 5.45 [Uncaptioned image] 6.75 [Uncaptioned image] 6.15 [Uncaptioned image] 5.59 [Uncaptioned image] 5.67 [Uncaptioned image] 5.22 [Uncaptioned image] 3.67 [Uncaptioned image] 5.79
Vehicle color statistics: [Uncaptioned image] White colors  [Uncaptioned image] Gray colors  [Uncaptioned image] Black colors  [Uncaptioned image] Other chromatic colors
TABLE I: Details of our new vehicle ReID dataset (Vehicle-3I). Color (%) summarizes the statistics of vehicle colors of each camera. The proportion of achromatic vehicles such as white, gray, and black is much higher than that of chromatic colored vehicles.
@@ V Datasets
Figure 6: Overall connections and vehicle flows in Vehicle-3I
Figure 7: Ratio of achromatic colors and chromatic colors in VeRi776 [14] and Vehicle-3I. White, gray and black colors are achromatic colors and the other colors are chromatic.
For various experiments, we used VeRi776 [14] vehicle re-identification (ReID) datasets, and Market-1501 [13] person ReID dataset. In addition, to validate the proposed methods in a real-world scenario, we newly constructed a Vehicle-3I vehicle ReID dataset.

### v2

@@ III-B Fusion Network
The baseline for the proposed ReID framework can be any appearance-based ReID method that estimates appearance similarities between images. Object images from each camera pair $(c_{i},c_{j})$ are represented as $\mathbf{I}^{m}_{i}$ and $\mathbf{I}^{n}_{j}$, where $m$ and $n$ are the image indices. The appearance-based ReID methods then estimate the visual similarity between two images as $S_{A}(\mathbf{I}^{m}_{i},\mathbf{I}^{n}_{j})$, which lies within the range $\left[0,1\right]$. The proposed framework does not rely only on typical appearance-based models.
To perform spatial-temporal ReID, Cho et al. [7] used only camera network topology to limit the gallery search range, effectively reducing the complexity of ReID. However, the spatial-temporal probability did not affect the final similarity. In contrast, previous studies [10, 33, 9] have merged both probabilities (i.e., appearance and spatial-temporal) with the same importance to obtain the joint probability. However, they neglected two key points. First, the domains of each probability are not the same. Second, both appearance and spatial-temporal probabilities can be imperfect. Hence, simply merging these probabilities is unreasonable.
In this study, we optimally combined visual similarities $S_{A}(\mathbf{I}^{m}_{i},\mathbf{I}^{n}_{j})$ and estimated spatial-temporal probability distributions $p_{ij}\left(\tau\right)$ through a fusion network named FusionNet. The input vector of the network for two images $(\mathbf{I}^{m}_{i},\mathbf{I}^{n}_{j})$ in the camera pair $(i,j)$ can be represented by
@@
where $S_{A}$ is the appearance similarity, and $S_{T}$ is the spatial-temporal vector. The $S_{T}$ vector between the images is defined by
@@
the notation $\tau^{mn}_{ij}$ represents the time difference between two images $\mathbf{I}^{m}_{i}$ and $\mathbf{I}^{n}_{j}$. The parameter $W$ denotes the size of a time window. The range of the $S_{T}$ vector depends on the value of $W$ and is determined around the distributions of $p_{ij}(\tau^{mn}_{ij})$. For example, when $W=0$, $S_{T}$ becomes a scalar value given by $S_{T}(\mathbf{I}^{m}_{i},\mathbf{I}^{n}_{j})=p_{ij}(\tau^{mn}_{ij})$. If $W>0$, the $S_{T}$ vector has a dimension of $12W+$ 1. The value of $W$ can be adjusted to determine the amount of spatial-temporal information provided to FusionNet. Consequently, the dimension of the input vector $\mathbf{x}^{mn}_{ij}$ for FusionNet is $22W+$ 2.
(a) One-to-all ReID without causality
(b) ID-to-ID ReID considering causality
Figure 4: Comparisons of ReID methodologies: Circles and boxes denote cameras and appearances. The number inside each box represents the appearance ID. Boxes filled with olive color indicate true positives concerning the query. Boxes with dotted lines indicate appearances excluded from the gallery. (a) One-to-all ReID using only a single appearance for the query performs many redundant and duplicated comparisons, even comparing different objects within the same camera.
@@
The FusionNet was designed based on a simple multi-layer perception. Empirical findings revealed that FusionNet does not require a sophisticated deep neural network structure to estimate the final similarity. The network consists of one hidden layer with several nodes and a one-dimensional output layer, as illustrated in Fig. 2(a). For the activation function, we used the Rectified Linear Unit (ReLU) for nodes in the hidden layer and the sigmoid function for the output node. Consequently, the final output of FusionNet, $S_{F}(\mathbf{I}^{m}_{i},\mathbf{I}^{n}_{j})$, is within the range $\left[0,1\right]$. To train the network, we optimized the binary cross-entropy loss, defined as follows:
@@ IV Re-identification in
Previous studies have evaluated their methods using a fixed and refined gallery set. They selected a query image from the gallery and compared it with all other images in the gallery, which is a one-to-all single-shot comparison. This evaluation process is repeated until all queries are compared, and then all the results are averaged to compute the final ReID performance. Note that the gallery set contains multiple appearances for a single identity, leading to many duplicated comparisons as shown in Fig. 4a. These laboratory evaluation settings are unsuitable for real-world scenarios, where multiple appearances of the object do not need to be evaluated individually.
Additional issues beyond laboratory settings must be considered in real-world surveillance scenarios. First, each identity can have multiple appearances in the video, and these appearances should be evaluated in an ID-to-ID manner, as explained in Fig. 4b. Second, the query and gallery sets are not fixed throughout all test processes; instead, they can change according to the target objects and matching causality. We propose CIM based on our previous work [12] to address these challenges in real-world scenarios. CIM includes three different methods (Sections IV-A– IV-C) for effective ReID in real-world scenarios. The overall framework of CIM is illustrated in Fig. 2(b).
@@ IV-A Appearance management for each object
As mentioned above, each object has multiple appearances along its trajectory in the sequence. Assuming that a camera captures multiple frames per second, we can collect various images, which is advantageous for ReID. However, this also requires a significant amount of computation. We observed that many of the appearances are unreliable and redundant in many cases. For example, when one object occludes another, the appearances of different objects can become very similar, resulting in unreliable appearances. Additionally, when an object remains in the same location and poses for a long period, its multiple appearances become identical, leading to redundancy. We define the bounding box of the $m$-th object image $\mathbf{I}^{m}$ as $\text{bbox}(\mathbf{I}^{m})=\left[x^{m},y^{m},w^{m},h^{m}\right]$. We omit the camera indices $i$ and $j$ at this step for convenience. In the video, the $m$-th object contains a set of appearances as
@@
where $N_{m}$ is the total number of appearances in $m$-th object.
@@
Figure 5: Feature distributions of objects. Example image patches reflect their actual sizes.
Inspired by [34], we perform appearance management for each object. To detect occluded unreliable objects, the occlusion ratio between two different objects is calculated based on Intersection over Union (IoU) as $\text{IoU}\left(\mathbf{I}^{m},\mathbf{I}^{n}\right)=\frac{\text{bbox}(\mathbf{I}^{m})\cap\text{bbox}(\mathbf{I}^{n})}{\text{bbox}(\mathbf{I}^{m})\cup\text{bbox}(\mathbf{I}^{n})}$. In general, surveillance cameras are installed at a high position from the ground, looking down at an angle. This means that an overlapping object with a larger $y$-value is occluding another object. Therefore, we designed the occlusion ratio for the $m$-th object as
@@
where $m\neq n$. When multiple bounding boxes occlude $\mathbf{I}^{m}$, the largest value will be used as the final occlusion ratio.
Next, we measure a self-overlapping ratio to prevent duplicate appearances from being included in the set $\mathcal{A}^{m}$. For simplicity, let $\mathbf{\hat{I}}^{m}\in\mathcal{A}^{m}$ represent the last appearance in the set $\mathcal{A}^{m}$. Then, the self-overlapping ratio is calculated using $\text{IoU}\left(\mathbf{I}^{m},\mathbf{\hat{I}}^{m}\right)$. Finally, we merge the two ratios to define the appearance unreliability score of $\mathbf{I}^{m}$ by
@@
Figure 5 illustrates the feature distributions of four different objects. Due to occlusions, these objects are not separated in the feature space without the proposed appearance management. Furthermore, each object has similar and redundant appearances when standing at the traffic signal. After performing appearance management, the unreliable appearances were appropriately removed, preserving the original distributions with a much smaller number of appearances.
@@ IV-B Top-$k$ Multi-shot Matching
After appearance management, objects contain more reliable appearances in their appearance sets. To compare two different appearance sets, many traditional multi-shot matching ReID methods have tried maximum (max) and average (avg.) aggregation approaches. However, the ‘max’ approach, which selects only the maximum similarity value from multiple appearance comparisons, can produce highly biased final results. On the other hand, the ‘avg.’ approach, which averages all comparison results, can result in overly smoothed final outcomes.
To mitigate these risks, we perform Top-$k$ multi-shot matching to compare two different appearance sets. Given two appearance sets $\mathcal{A}^{m},\mathcal{A}^{n}$ with cardinalities $|\mathcal{A}^{m}|=N_{m}$, $|\mathcal{A}^{n}|=N_{n}$, respectively, there are $N_{m}\times N_{n}$ possible appearance comparisons. Using the FusionNet described in Section III-B, the appearance similarities between the two sets are represented by
@@
Sort $S_{F}\left(\mathcal{A}^{m},\mathcal{A}^{n}\right)$ in descending order and obtain
@@
Finally, the top-$k$ multi-shot matching is calculated by
@@
By adjusting the value $k$, we can control how many results are used in the multi-shot appearance comparison. We empirically set $k$ to around $5$, which resulted in improved ReID performances, as shown in Table. VI. The proposed top-$k$ approach is quite simple but effective. It does not require complex structures for handling multi-shot comparisons like video-based ReID methods [35, 36] but can utilize single-shot baselines. In addition, it implements other approaches by setting $k=1$ for the ‘max’ approach and $k=N_{m}\times N_{n}$ for the ‘avg.’ approach.
@@ IV-C Causal Identity Matching (CIM)
In Section III-A, the camera network topology was built as $\mathbf{G}=(\mathbf{V},\mathbf{E})$, and FusionNet was trained based on this topology. The topology includes transition time distributions $p_{ij}$ between all possible camera pairs, regardless of their connection strengths, allowing FusionNet to learn these strengths to determine the similarity $S_{F}$. In real-world scenarios, we focus more on the target object (i.e., query) and its causality in the camera networks $\mathbf{G}$. The causality indicates where the target object is currently and where it is likely to move next based on the topology. In this setting, the query and gallery sets should not be fixed or pre-defined but should dynamically change according to the ReID results at each step.
To consider aforementioned conditions and perform causal identity matching (CIM), we first define the camera adjacent matrix by $\mathbf{C}\in\{c_{ij}|1\leq i\leq N_{cam},1\leq j\leq N_{cam}\}$, where $i,j$ are camera indexes, and $c_{ij}$ denotes a binary adjacent flag between cameras. When cameras $c_{i}$ and $c_{j}$ are adjacent, the flag is set to $c_{ij}=1$; otherwise, it is set to $c_{ij}=0$. All diagonal elements of $\mathbf{C}$ are set to $c_{ii}=0$. Given multiple object paths, we can simply find adjacent cameras and their overall connections. For example, assuming there are multiple object paths such as $\left(c_{1}\rightarrow c_{2}\rightarrow c_{3}\right)$, $\left(c_{2}\rightarrow c_{3}\right)$, $\left(c_{2}\rightarrow c_{4}\right)$, and $\left(c_{4}\rightarrow c_{2}\right)$; the camera adjacent matrix is defined by
@@
Note that the matrix can be asymmetric when considering a bidirectional camera network topology. More reliable adjacent connections can be estimated by voting on sufficient numbers of paths. Additionally, unnecessary paths can be removed through post-processing algorithms like transitive reduction [37] [^1] When there is a new path $\left(c_{1}\rightarrow c_{3}\right)$, it does not assign $c_{13}=1$, immediately. Instead, the algorithm tries to infer if there is an intermediate vertex $c_{2}$ between $c_{1}$ and $c_{3}$ based on numerous training paths.. Excluding redundant paths helps to improve the overall efficiency of ReID.
@@
Figure 6: Overall connections and vehicle flows in the Vehicle-3I
Next, we assign dynamic gallery sets for the target object. Suppose a target object (query) is detected by camera $c_{1}$. Then, the gallery set of the query comes from the cameras adjacent to $c_{1}$ by finding $(c_{1j}=1)$ in $\mathbf{C}$. Furthermore, the objects in the gallery are verified by referring to the corresponding transition time distributions $p_{1j}\left(t\right)$. Only objects within the camera $j$ that appear within a time $t$ satisfying $p_{1j}\left(t\right)>0$ are included in the gallery. Through this process, we can causally assign the gallery set using $\mathbf{C}$ and reduce many unnecessary matching candidates based on $p_{ij}\left(t\right)$.
An object in the gallery with the highest similarity to the query is considered a successful match, and the query is merged with the re-identified object. This increases the diversity of the query’s appearance, making the subsequent ReID steps more advantageous. In other words, although a single camera can only capture fixed poses of the object, progressive query merging easily overcomes this limitation. The updated query is then used to progressively perform ReID on the following adjacent cameras, with the gallery adaptively determined by $\mathbf{C}$ and $p_{ij}$.
V Datasets
Figure 7: The ratio of achromatic to chromatic colors in the VeRi776 [14] and Vehicle-3I datasets is provided. White, gray, and black are categorized as achromatic colors, while all other colors are chromatic.
For various experiments, we used VeRi776 [14] vehicle re-identification (ReID) datasets, and Market-1501 [13] person ReID dataset. In addition, to validate the proposed methods in a real-world scenario, we newly constructed the Vehicle-3I vehicle ReID dataset.